OpenStudy (anonymous):

An ant looks up to the top of a tree at an angle of elevation of 40 degrees. If the ant is 30 feet from the base of the tree, what is the height of the tree, to the nearest tenth?

OpenStudy (0_0youscareme):

http://openstudy.com/study#/updates/539f0738e4b0eb8de56e5c68 read threw that

OpenStudy (0_0youscareme):

This is a right triangle problem using the tangent function. Tan(x) = Opposite (O) / Adjacent (A) x = 40 degrees A = 30 ft Solve for O O= Tan (x) * A O = Tan (40) * 30 Use your calculator or whatever method to get the tangent of 40 degrees (which is .84) O = .84 * 30 O = 25.2 Ft

OpenStudy (anonymous):